OpenStudy (anonymous):

More help plzzz Solve this system by elimination. Express your answer as an ordered pair. 2x - 5y = -27 2x - 3y = -13

OpenStudy (btaylor):

Elimination means that we have to get rid of one part. Multiply the top by -1: -2x+5y=27 2x - 3y=-13 ------------ 0 +2y = 14 --> y=7 2x-3(7)=-13 2x-21=13 2x=8 x=4

OpenStudy (anonymous):

thx dats all i need to know

OpenStudy (btaylor):

yw

OpenStudy (anonymous):

you mean you just need the answer? LOL

OpenStudy (anonymous):

lol yes

OpenStudy (anonymous):

@BTaylor You just wasted your time XD

OpenStudy (anonymous):

i was joking keep going

OpenStudy (anonymous):

not relly i gave him medal

OpenStudy (btaylor):

@sourwing i know but "give a man a fish, and you feed him for a day. teach him how to fish and you feed him for a lifetime."

OpenStudy (anonymous):

wow thats da best quote ever

OpenStudy (anonymous):

doesn't apply to this case XD